# Markov Decision Processes                                  Under External Temporal Processes

### v2

@@ 9. Discussion and Future Directions
The theorem proposed is general and relies on learning functions in an augmented state space that blows up in the presence of long-term dependencies due to the external events. For a practical algorithm, it is necessary to compress this augmented state space to something more manageable.
This can be done implicitly by using function approximators such as recurrent neural networks or LSTMs, or explicitly by learning compressed state representations using mutual information bottlenecks etc. While these are practical methods, it is also an open problem on how to theoretically analyze such a learnt compressed state representation.
@@
The sample complexity analysis in Section 7 hinges on generalizing sample error of policy evaluation to an expected error over the entire state space. Such generalization generally requires assumptions on the distribution over states.
In the theorem stated by us, the assumption was on the rate of convergence of the Markov chain induced from the augmented MDP by the policy under evaluation, which is common in the reinforcement learning literature. However, in our case, the specific structure of the state space, i.e, separation of the augmented state space into states and events with different characteristics, may lead itself to a different types of assumptions. For example, the recurrence and aperiodicity of the Markov chain may not be a good assumption, but a stationarity assumption on just the event sequence might be justified.
@@
The policy improvement step involves optimizing over the action space based on the expected one-step reward and the value function at the next time step, which requires knowledge of the model. Thus, in the case of an unknown transition and reward model, it is important to investigate the sample complexity of learning the models from samples.
Appendix A Proofs of lemmas
A.1. Proof of Lemma  5.2
We adapt the proof of Theorem 1 from Xiao et al. (2019).
We wish to study the value function of the policy in two different MDPs $\mathcal{M}_{X}$ and $\mathcal{M}^{(T)}$. For convenience, without any loss in generality, we can extend the state space of $\mathcal{M}^{(T)}$ to that of $\mathcal{M}_{X}$ by including the information of every past event in the state, without changing the transition function.
We define another class of auxiliary MDPs $\mathcal{M}^{(T)}_{h}$, $h\in 0\cup\mathbb{N}$, in which when starting from initial (augmented) state $\bar{s}\in\bar{S}$, the transitions follow the transition kernel of $\mathcal{M}_{X}$ for $h$ time steps and then the transition kernel of $\mathcal{M}^{(T)}$ for all transitions thereafter. The value functions induced by $\pi$ in $\mathcal{M}_{X}$ and $\mathcal{M}^{(T)}$ can be related using these new intermediate MDPs, which interpolate between those two environments.
Since only one policy is under discussion, the $\pi$ in $V^{\pi}$ is omitted for the rest of this proof for simplicity of notation. $V$ denotes the value function of the policy in $\mathcal{M}_{X}$, $V^{(T)}$ denotes its value in $\mathcal{M}^{(T)}$, and $V_{h}^{(T)}$ in $\mathcal{M}^{(T)}_{h}$.
The following relations follow from the definitions of the auxiliary and intermediate MDPs and the boundedness of $r$:
@@ Appendix A Proofs of lemmas
That is, the entire difference can be characterized in terms of differences in $V_{h}^{(T)}$ and $V_{h+1}^{(T)}$ for each $h$. Now for any $s\in\bar{S}$,
$V_{h}^{(T)}(s)$ $=\sum_{t=0}^{h-1}\gamma^{t}\mathsf{E}_{\begin{subarray}{c}s_{t}\sim p_{t}(.|s)\\ a_{t}\sim\pi(.|s_{t})\\ s_{t+1}\sim p(.|s_{t},a_{t})\end{subarray}}\left[r(s_{t},a_{t},s_{t+1})\right]$
$\quad+\gamma^{h}\mathsf{E}_{\begin{subarray}{c}s_{h}\sim p_{h}(.|s)\\ a_{h}\sim\pi(.|s_{h})\\ s_{t+1}\sim p^{(T)}(.|s_{h},a_{h})\end{subarray}}\left[r(s_{h},a_{h},s_{h+1})\right]$
$\quad+\sum_{t=h+1}^{\infty}\gamma^{t}\mathsf{E}_{\begin{subarray}{c}s_{t}\sim p_{t-h}^{(T)}\circ p_{h}(.|s)\\ a_{t}\sim\pi(.|s_{t})\\ s_{t+1}\sim p^{(T)}(.|s_{t},a_{t})\end{subarray}}\left[r(s_{t},a_{t},s_{t+1})\right],$
where $p_{t}(.|s)$ denotes a transition of $t$ steps in the environment $\mathcal{M}_{X}$ starting from $s$, a superscript (T) denotes the corresponding object in $\mathcal{M}^{(T)}$, and the policy $\pi$ is implicit whenever it is not explicitly mentioned in any expression.
@@
$V_{h}^{(T)}(s)=\sum_{t=0}^{h-1}\gamma^{t}\mathsf{E}_{\begin{subarray}{c}s_{t}\sim p_{t}(.|s)\\ a_{t}\sim\pi(.|s_{t})\\ s_{t+1}\sim p(.|s_{t},a_{t})\end{subarray}}\left[r(s_{t},a_{t},s_{t+1})\right]+\gamma^{h}\mathsf{E}_{\begin{subarray}{c}s_{h}\sim p_{h}(.|s)\end{subarray}}V^{(T)}(s_{h}).$
@@
$V_{h+1}^{(T)}(s)-V_{h}^{(T)}(s)$ $=\gamma^{h}\mathsf{E}_{\begin{subarray}{c}s_{h}\sim p_{h}(.|s)\\ a_{h}\sim\pi(.|s_{h})\\ s_{h+1}\sim p(.|s_{h},a_{h})\end{subarray}}\left[r(s_{h},a_{h},s_{h+1})\right]$
$\qquad\qquad\qquad\qquad\qquad+\gamma^{h+1}\mathsf{E}_{\begin{subarray}{c}s_{h+1}\sim p_{h+1}(.|s)\end{subarray}}V^{(T)}(s_{h+1})$
$-\gamma^{h}\mathsf{E}_{\begin{subarray}{c}s_{h}\sim p_{h}(.|s)\\ a_{h}\sim\pi(.|s_{h})\\ s_{t+1}\sim p^{(T)}(.|s_{h},a_{h})\end{subarray}}\left[r(s_{h},a_{h},s_{h+1})\right]$
$\qquad\qquad\qquad\qquad\qquad+\gamma^{h+1}\mathsf{E}_{\begin{subarray}{c}s_{h+1}\sim p^{(T)}\circ p_{h}(.|s)\end{subarray}}V^{(T)}(s_{h+1})$
$=\gamma^{h}\mathsf{E}_{\begin{subarray}{c}s_{h}\sim p_{h}(.|s)\\ a_{h}\sim\pi(.|s_{h})\end{subarray}}\bigg{[}\mathsf{E}_{s_{h+1}\sim p(.|s_{h},a_{h})}r(s_{h},a_{h},s_{h+1})$
$\qquad\qquad\qquad\qquad\qquad-\mathsf{E}_{s_{h+1}\sim p^{(T)}(.|s_{h},a_{h})}r(s_{h},a_{h},s_{h+1})\bigg{]}$
$+\gamma^{h+1}\mathsf{E}_{\begin{subarray}{c}s_{h}\sim p_{h}(.|s)\\ a_{h}\sim\pi(.|s_{h},a_{h})\end{subarray}}\bigg{[}\mathsf{E}_{s_{h+1}\sim p(.|s_{h},a_{h})}V^{(T)}(s_{h+1})$
$\qquad\qquad\qquad\qquad\qquad-\mathsf{E}_{s_{h+1}\sim p^{(T)}(.|s_{h},a_{h})}V^{(T)}(s_{h+1})\bigg{]}.$
In the above equations, $s,s_{t}$ etc., denote the augmented state, which consists of the state and event mark history till then. Since the current state and the current event mark are independent and depend only on the previous events, the transition kernel can be factored as two independent distributions over $S$ and $\mathcal{X}$. So, replacing the augmented state $s\in\bar{S}$ with the explicit representation $(s,H)$, where $s\in S$ and $H\in\mathcal{X}^{\infty}$ gives, for any bounded measurable function $f$ on $\mathbb{K}\times\mathcal{X}^{\infty}$,
@@
$=\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})-\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X(T)}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})$
$=\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})-\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})$
$+\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})-\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X(T)}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})$
$\leq\mathsf{E}_{\begin{subarray}{c}X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}\left[\left|\left|f(.,X_{t+1},H_{t})\right|\right|_{\infty}\mathsf{TV}\left(Q_{H_{t}}(.|s_{t},a_{t}),Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\right)\right]$
$+\mathsf{E}_{s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})}\left[\left|\left|f(s_{t+1},.,H_{t})\right|\right|_{\infty}\mathsf{TV}\left(Q^{X}_{H_{t}}(.|s_{t},a_{t}),Q^{X(T)}_{H_{t}}(.|s_{t},a_{t})\right)\right]$
$\leq||f||_{\infty}\left(\sum_{t=T+1}^{\infty}M_{t}+\sum_{t=T+1}^{\infty}N_{t}\right).$
The second to last inequality is a result of Total Variation Distance being an integral probability metric (Sriperumbudur et al., 2009). Using this expression in the previous equation gives
$V_{h+1}^{(T)}(s)-V_{h}^{(T)}(s)$ $\leq\gamma^{h}||r||_{\infty}\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
$\qquad\qquad+\gamma^{h+1}||V^{(T)}||_{\infty}\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
$\leq\gamma^{h}\left(||r||_{\infty}+\gamma||V^{(T)}||_{\infty}\right)\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right).$
@@
$||V-V^{(T)}||_{\infty}$ $\leq\sum_{h=1}^{\infty}\gamma^{h}\left(||r||_{\infty}+\gamma||V^{(T)}||_{\infty}\right)\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
$=\frac{1}{1-\gamma}\left(||r||_{\infty}+\gamma||V^{(T)}||_{\infty}\right)\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right).$
A.2. Proof of Lemma 6.1
@@
Taking the difference between these two expressions, and adding and subtracting the additional term
@@
$\leq||r||_{\infty}\left(\sum_{t=T+1}^{\infty}M_{t}\right)+||V^{\pi}||_{\infty}\left(\sum_{t=T+1}^{\infty}M_{t}+N_{t}\right)$
$\qquad\qquad+\gamma\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T}}(.|s,a)\\ x^{\prime}\sim Q_{x_{0:T}}(.)\\ a=\pi((s,x_{0:T}))\end{subarray}}\left|V^{\pi}((s^{\prime},x^{\prime},x_{0:\infty}))-V^{\pi}((s^{\prime},x^{\prime},x_{0:T}))\right|.$
@@
Since each repetition adds a $\gamma$ factor to the term, which itself is bounded by $\frac{1}{1-\gamma}$ because of bounded rewards, the resulting series converges, giving us the following bound:
@@
$\leq\sum_{t=T+1}^{\infty}\gamma^{t-T-1}\left[||r||_{\infty}\left(\sum_{t^{\prime}=t}^{\infty}M_{t^{\prime}}\right)+||V||_{\infty}\left(\sum_{t^{\prime}=t}^{\infty}M_{t^{\prime}}+N_{t^{\prime}}\right)\right]$
$\leq\left(1+\frac{\gamma}{1-\gamma}\right)\sum_{t=T+1}^{\infty}\sum_{t^{\prime}=t}^{\infty}\gamma^{t}(M_{t^{\prime}}+N_{t^{\prime}})$
$\leq\frac{1}{1-\gamma}\sum_{t^{\prime}=T+1}^{\infty}\sum_{t=0}^{T-t^{\prime}-1}\gamma^{t}(M_{t^{\prime}}+N_{t^{\prime}})$
$=\frac{1}{1-\gamma}\sum_{t^{\prime}=T+1}^{\infty}\frac{1-\gamma^{t^{\prime}-T}}{1-\gamma}(M_{t^{\prime}}+N_{t^{\prime}})\leq\frac{1}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right).$
Appendix B Details of Theorem 7.1
@@
A Markov chain $\mathcal{M}=\left(X_{t}\right)_{t\geq 1}$ is said to be exponentially $\beta$-mixing with parameters $\bar{\beta},b,\kappa$ if its mixing coefficients
@@
When such an exponentially mixing Markov chain is also ergodic and and aperiodic with stationary distribution $\rho$, for any initial distribution $\lambda$, there is a bound on the following total variation: (Definition 21 of Lazaric et al. (2012))
$\left|\left|\int_{\mathcal{X}}\lambda(dx)\mathsf{P}\left(.|x\right)-\rho(.)\right|\right|_{\mathsf{TV}}\leq\bar{\beta}\exp\left(-bi^{\kappa}\right).$

### v3

@@ 9. Discussion and Future Directions
This work sets the stage for understanding problems related to RL under the influence of external temporal events and provides several theoretical insights. The results are established in a very general setting and rely on learning functions in an augmented state space that is exponentially bigger in the presence of long-term dependencies due to external events. To mitigate this, one needs to develop provable approximate methods by reducing the size of the state space. Due to the ”factorization” of the environment into Markovian states and non-Markovian events that impact these states, compressing the histories of events into latent features that are independent of the states may provide a strategic advantage. This approach could facilitate the learning of representations of the expanded states by developing a distinct event distribution model that can be scaled up, as it does not rely on the agent, alongside a separate state transition model that necessitates agent interaction with the environment. A compressed representation of events, derived from extensive event data, can subsequently be utilized to learn a feasible state transition kernel.
This can be achieved either implicitly through function approximators such as recurrent neural networks (or LSTMs), or explicitly by learning compressed state representations using mutual information bottlenecks, etc. While these are practical methods, the theoretical analysis of such learned compressed state representations remains an open problem.
@@
The analysis of sample complexity in Section 6 hinges on generalizing sample error of policy evaluation to an expected error over the entire state space. Such generalization generally requires assumptions on the distribution over states.
In our result, the assumption pertains to the rate of convergence of the Markov chain induced from the augmented MDP by the policy under evaluation, a notion frequently encountered in reinforcement learning literature. However, in our case, the specific structure of the state space–specifically, the separation of the augmented state space into states and events with different characteristics–may necessitate alternate assumptions. For instance, while the recurrence and aperiodicity of the Markov chain may not hold, it may be reasonable to assume stationarity solely for the event sequence.
@@
The policy improvement step involves optimizing over the action space based on the expected one-step reward and the value function at the next time step, which requires knowledge of the model. Thus, in the case of an unknown transition and reward model, it is important to investigate the sample complexity of learning the models from samples and how this affects the performance of the resultant policy.
Appendix A Notation
Table 1. Table of notation
$\mathcal{M}$ Stationary MDP
$S$ State space
$A$ Action space
$Q$ Transition probability kernel
$r$ Reward function
$r(s,a)$ Reward for taking action $a$ at state $s$
$r(s,a,s^{\prime})$ Reward if taking action $a$ at state $s$ led to next state $s^{\prime}$
$\gamma$ Discount factor
$\pi$ Policy
$V^{\pi}$ Value function of policy $\pi$
$\mathcal{T}$ Optimal Bellman operator
$\pi^{*}$ Optimal policy
$V^{*}$ Value function of $\pi^{*}$
$\mathsf{TV}(.\;,\;.)$ Total Variation distance between two probability distributions
$\mathcal{X}$ Space of event marks
$\mathcal{M}_{X}$ Augmented MDP
$\mathcal{M}^{(T)}_{X}$ Approximate augmented MDP with finite history of $T+1$ events
$H$ History of external events till time $t$
$Q_{H}$ Transition kernel of the MDP due to influence of event history $H$
$Q^{X}_{H}$ Distribution of the mark of the event process due to history $H$
$\bar{S}$ Augmented state space
$\bar{s}$ State in $\bar{S}$
$x$ Event mark, an element of $\mathcal{X}$
$x_{0:t}$ Events that occurred withing the past $t$ time steps
$x_{t:\infty}$ Events older than $t$ time steps
$\left(x_{0:t},x^{\prime}_{t+1:\infty}\right)$ Event history obtained by concatenating past few events of $x$
with older events from $x^{\prime}$
$\varphi_{i}$ Basis functions
$\phi$ Feature function consisting of all basis functions
$\Phi$ Matrix containing feature vectors
$d$ Dimensionality of feature representation, i.e., number of basis functions
$\hat{P}$ Empirical transition matrix
$\hat{V}$ Value function estimated using pathwise LSTD
$\widetilde{V}$ $\hat{V}$ clipped at $\frac{1}{1-\gamma}$
$\nu$ Lower bound on the eigenvalues of $\frac{1}{N}\Phi^{T}\Phi$
Appendix B Proofs of lemmas
B.1. Proof of Lemma  4.2
The aim is to study the value function of the policy in two MDPs $\mathcal{M}_{X}$ and $\mathcal{M}^{(T)}_{X}$, where $\mathcal{M}_{X}$ is the non-stationary MDP with augmented states consisting of the actual state and the entire history of events till then, and $\mathcal{M}^{(T)}_{X}$ is an approximate MDP where only events in the past $T$ time steps affect the state transition and event distribution. For convenience, without any loss in generality, one can extend the state space of $\mathcal{M}^{(T)}_{X}$ to that of $\mathcal{M}_{X}$ by including the information of every past event in the state, without changing the transition function.
Now, We define a class of auxiliary MDPs $\mathcal{M}^{(T)}_{h}$, $h\in\left\{0,1,2,\dots\right\}$, in which when starting from initial (augmented) state $\bar{s}\in\bar{S}$, the transitions follow the transition kernel of $\mathcal{M}_{X}$ for $h$ time steps and then the transition kernel of $\mathcal{M}^{(T)}_{X}$ for all transitions thereafter. The value functions induced by $\pi$ in $\mathcal{M}_{X}$ and $\mathcal{M}^{(T)}_{X}$ can be related using these new intermediate MDPs, which interpolate between those two environments.
Since only a single policy is under discussion, the $\pi$ in $V^{\pi}$ is omitted for the rest of this proof for simplicity of notation. $V$ denotes the value function of the policy in $\mathcal{M}_{X}$, $V^{(T)}$ denotes its value in $\mathcal{M}^{(T)}_{X}$, and $V_{h}^{(T)}$ in $\mathcal{M}^{(T)}_{h}$.
The following relations follow from the definitions of the auxiliary and intermediate MDPs and the boundedness of $r$.
@@ Appendix B Proofs of lemmas
That is, the difference between $V$ and $V^{(T)}$ can be characterized in terms of differences in $V_{h}^{(T)}$ and $V_{h+1}^{(T)}$ for each $h\in\{0,1,2,\dots\}$. Now, for any augmented state $\bar{s}\in\bar{S}$,
$V_{h}^{(T)}(\bar{s})$ $=\sum_{t=0}^{h-1}\gamma^{t}\mathsf{E}_{\begin{subarray}{c}\bar{s}_{t}\sim p_{t}(.|\bar{s})\\ a_{t}\sim\pi(.|\bar{s}_{t})\\ \bar{s}_{t+1}\sim p(.|\bar{s}_{t},a_{t})\end{subarray}}\left[r(\bar{s}_{t},a_{t},\bar{s}_{t+1})\right]+\gamma^{h}\mathsf{E}_{\begin{subarray}{c}\bar{s}_{h}\sim p_{h}(.|\bar{s})\\ a_{h}\sim\pi(.|\bar{s}_{h})\\ \bar{s}_{t+1}\sim p^{(T)}(.|\bar{s}_{h},a_{h})\end{subarray}}\left[r(\bar{s}_{h},a_{h},\bar{s}_{h+1})\right]$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\sum_{t=h+1}^{\infty}\gamma^{t}\mathsf{E}_{\begin{subarray}{c}\bar{s}_{t}\sim p_{t-h}^{(T)}\circ p_{h}(.|\bar{s})\\ a_{t}\sim\pi(.|\bar{s}_{t})\\ \bar{s}_{t+1}\sim p^{(T)}(.|\bar{s}_{t},a_{t})\end{subarray}}\left[r(\bar{s}_{t},a_{t},\bar{s}_{t+1})\right],$
where $p_{t}(.|\bar{s})$ denotes a transition of $t$ steps in the environment $\mathcal{M}_{X}$ starting from $\bar{s}$, a superscript (T) denotes the corresponding object in $\mathcal{M}^{(T)}_{X}$, and the policy $\pi$ is implicit whenever it is not explicitly mentioned in any expression. The reward function does not depend on the external events, so $r(\bar{s}_{t},a_{t},\bar{s}_{t+1)}$ is just $r(s_{t},a_{t},s_{t+1})$, where $\bar{s}=(s,x)$ and $s$ is the actual state.
@@
$V_{h}^{(T)}(\bar{s})=\sum_{t=0}^{h-1}\gamma^{t}\mathsf{E}_{\begin{subarray}{c}\bar{s}_{t}\sim p_{t}(.|\bar{s})\\ a_{t}\sim\pi(.|\bar{s}_{t})\\ \bar{s}_{t+1}\sim p(.|\bar{s}_{t},a_{t})\end{subarray}}\left[r(\bar{s}_{t},a_{t},\bar{s}_{t+1})\right]+\gamma^{h}\mathsf{E}_{\begin{subarray}{c}\bar{s}_{h}\sim p_{h}(.|\bar{s})\end{subarray}}V^{(T)}(\bar{s}_{h}).$
@@
$V_{h+1}^{(T)}(\bar{s})-V_{h}^{(T)}(\bar{s})$ $=\gamma^{h}\mathsf{E}_{\begin{subarray}{c}\bar{s}_{h}\sim p_{h}(.|\bar{s})\\ a_{h}\sim\pi(.|\bar{s}_{h})\\ \bar{s}_{h+1}\sim p(.|\bar{s}_{h},a_{h})\end{subarray}}\left[r(\bar{s}_{h},a_{h},\bar{s}_{h+1})\right]$
$\qquad\qquad\qquad\qquad\qquad+\gamma^{h+1}\mathsf{E}_{\begin{subarray}{c}\bar{s}_{h+1}\sim p_{h+1}(.|\bar{s})\end{subarray}}V^{(T)}(\bar{s}_{h+1})$
$\qquad\qquad-\gamma^{h}\mathsf{E}_{\begin{subarray}{c}\bar{s}_{h}\sim p_{h}(.|\bar{s})\\ a_{h}\sim\pi(.|\bar{s}_{h})\\ \bar{s}_{t+1}\sim p^{(T)}(.|\bar{s}_{h},a_{h})\end{subarray}}\left[r(\bar{s}_{h},a_{h},\bar{s}_{h+1})\right]$
$\qquad\qquad\qquad\qquad\qquad+\gamma^{h+1}\mathsf{E}_{\begin{subarray}{c}\bar{s}_{h+1}\sim p^{(T)}\circ p_{h}(.|\bar{s})\end{subarray}}V^{(T)}(\bar{s}_{h+1})$
$=\gamma^{h}\mathsf{E}_{\begin{subarray}{c}\bar{s}_{h}\sim p_{h}(.|\bar{s})\\ a_{h}\sim\pi(.|\bar{s}_{h})\end{subarray}}\bigg{[}\mathsf{E}_{\bar{s}_{h+1}\sim p(.|\bar{s}_{h},a_{h})}r(\bar{s}_{h},a_{h},\bar{s}_{h+1})$
$\qquad\qquad\qquad\qquad\qquad-\mathsf{E}_{\bar{s}_{h+1}\sim p^{(T)}(.|\bar{s}_{h},a_{h})}r(\bar{s}_{h},a_{h},\bar{s}_{h+1})\bigg{]}$
$\qquad\qquad+\gamma^{h+1}\mathsf{E}_{\begin{subarray}{c}\bar{s}_{h}\sim p_{h}(.|\bar{s})\\ a_{h}\sim\pi(.|\bar{s}_{h},a_{h})\end{subarray}}\bigg{[}\mathsf{E}_{\bar{s}_{h+1}\sim p(.|\bar{s}_{h},a_{h})}V^{(T)}(\bar{s}_{h+1})$
$\qquad\qquad\qquad\qquad\qquad-\mathsf{E}_{\bar{s}_{h+1}\sim p^{(T)}(.|\bar{s}_{h},a_{h})}V^{(T)}(\bar{s}_{h+1})\bigg{]}.$
Since the current state and the current event mark are independent and depend only on the previous events, the transition kernel can be factored as two independent distributions over $S$ and $\mathcal{X}$. Hence, replacing the augmented state $\bar{s}\in\bar{S}$ with the explicit representation $(s,H)$, where $s\in S$ and $H\in\mathcal{X}^{\infty}$ gives, for any bounded measurable function $f$ on $S\times\mathcal{X}^{\infty}$,
@@
$\qquad=\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})-\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X(T)}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})$
$\qquad=\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})-\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})$
$\qquad\qquad\qquad+\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})-\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X(T)}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})$
$\qquad\leq\mathsf{E}_{\begin{subarray}{c}X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}\left[\left\|f(.,X_{t+1},H_{t})\right\|_{\infty}\mathsf{TV}\left(Q_{H_{t}}(.|s_{t},a_{t}),Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\right)\right]$
$\qquad\qquad\qquad+\mathsf{E}_{s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})}\left[\left\|f(s_{t+1},.,H_{t})\right\|_{\infty}\mathsf{TV}\left(Q^{X}_{H_{t}}(.|s_{t},a_{t}),Q^{X(T)}_{H_{t}}(.|s_{t},a_{t})\right)\right]$
$\qquad\leq\left\|f\right\|_{\infty}\left(\sum_{t=T+1}^{\infty}M_{t}+\sum_{t=T+1}^{\infty}N_{t}\right).$
The second to last inequality is a result of Total Variation Distance being an integral probability metric (Sriperumbudur et al., 2012). Using this expression in the previous equation gives
$V_{h+1}^{(T)}(\bar{s})-V_{h}^{(T)}(\bar{s})$ $\leq\gamma^{h}\left\|r\right\|_{\infty}\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
$\qquad\qquad+\gamma^{h+1}\left\|V^{(T)}\right\|_{\infty}\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
$\leq\gamma^{h}\left(\left\|r\right\|_{\infty}+\gamma\left\|V^{(T)}\right\|_{\infty}\right)\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right).$
@@
$\left\|V-V^{(T)}\right\|_{\infty}$ $\leq\sum_{h=1}^{\infty}\gamma^{h}\left(\left\|r\right\|_{\infty}+\gamma\left\|V^{(T)}\right\|_{\infty}\right)\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
$=\frac{1}{1-\gamma}\left(\left\|r\right\|_{\infty}+\gamma\left\|V^{(T)}\right\|_{\infty}\right)\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right).$
B.2. Proof of Lemma 5.1
Let $\pi$ be a policy that depends only on the current state and the past $T$ events.
@@
Taking the difference between these two expressions and adding and subtracting the additional term
@@
$\qquad\qquad\qquad\leq\left\|r\right\|_{\infty}\left(\sum_{t=T+1}^{\infty}M_{t}\right)+\left\|V^{\pi}\right\|_{\infty}\left(\sum_{t=T+1}^{\infty}M_{t}+N_{t}\right)$
$\qquad\qquad\qquad\qquad\qquad+\gamma\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T}}(.|s,a)\\ x^{\prime}\sim Q_{x_{0:T}}(.)\\ a=\pi((s,x_{0:T}))\end{subarray}}\left|V^{\pi}((s^{\prime},x^{\prime},x_{0:\infty}))-V^{\pi}((s^{\prime},x^{\prime},x_{0:T}))\right|.$
@@
Since each repetition adds a $\gamma$ factor to the term, which itself is bounded by $\frac{1}{1-\gamma}$ because of bounded rewards, the resulting series converges, giving us the following bound.
@@
$\qquad\qquad\qquad\qquad\leq\sum_{t=T+1}^{\infty}\gamma^{t-T-1}\left[\left\|r\right\|_{\infty}\left(\sum_{t^{\prime}=t}^{\infty}M_{t^{\prime}}\right)+\left\|V\right\|_{\infty}\left(\sum_{t^{\prime}=t}^{\infty}M_{t^{\prime}}+N_{t^{\prime}}\right)\right]$
$\qquad\qquad\qquad\qquad\leq\left(1+\frac{\gamma}{1-\gamma}\right)\sum_{t=T+1}^{\infty}\sum_{t^{\prime}=t}^{\infty}\gamma^{t}(M_{t^{\prime}}+N_{t^{\prime}})$
$\qquad\qquad\qquad\qquad\leq\frac{1}{1-\gamma}\sum_{t^{\prime}=T+1}^{\infty}\sum_{t=0}^{T-t^{\prime}-1}\gamma^{t}(M_{t^{\prime}}+N_{t^{\prime}})$
$\qquad\qquad\qquad\qquad=\frac{1}{1-\gamma}\sum_{t^{\prime}=T+1}^{\infty}\frac{1-\gamma^{t^{\prime}-T}}{1-\gamma}(M_{t^{\prime}}+N_{t^{\prime}})$
$\qquad\qquad\qquad\qquad\leq\frac{1}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right).$
Appendix C Details of Theorem 6.1
@@
A Markov chain $\left(X_{t}\right)_{t\geq 1}$ is said to be exponentially $\beta$-mixing with parameters $\bar{\beta},b,\kappa$ if its mixing coefficients
@@
When such an exponentially mixing Markov chain is also ergodic and and aperiodic with stationary distribution $\rho$, for any initial distribution $\lambda$, there is a bound on the following total variation: (Definition 21 in Lazaric et al. (2012))
$\left\|\int_{\mathcal{X}}\lambda(dx)\mathsf{P}\left(.|x\right)-\rho(.)\right\|_{\mathsf{TV}}\leq\bar{\beta}\exp\left(-bi^{\kappa}\right).$